OpenStudy (anonymous):

Solve the equation. will you please help me 78 = -2(m + 3) + m

OpenStudy (anonymous):

By distributive property 78=-2m-6+m 78=-2m+m-6 78=m-6 Adding 6 on both sides 78+6=m 84=m And the value of m is 84

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome